OpenStudy (anonymous):

Two polygons are congruent and the perimeter of the first polygon is 30 cm. If the sides of the second polygon are consecutive integers (x, x+1, x+2, x+3, etc.), what value of x makes the polygons into congruent quadrilaterals? Use x as the smallest side

OpenStudy (anonymous):

3 4 5 6

OpenStudy (callisto):

6

OpenStudy (callisto):

4x = 30-1-2-3; x=6

OpenStudy (anonymous):

Two polygons are congruent and the perimeter of the first polygon is 30 cm.? If the sides of the second polygon are consecutive integers (x, x+1, x+2, x+3, etc.), what value of x makes the polygons into congruent quadrilaterals? 4x+6=30 4x=24 x=6

OpenStudy (anonymous):

Answer: 6